4) If the reaction SO2Cl2 → SO2 + Cl2 is first-order reaction. At 230oC, k = 0.125 hr-1.

  1. If the initial concentration of SO2Cl2 is 0.250 M, what is the concentration after 7.6 hours?
  2. How long will it take for the concentration of SO2Cl2 to drop to 0.0651 M?
  3. What is the half-life for this first order reaction?
Answer #1

A. ln(A°/A) = rate constant * time

= ln ( 0.250 / A) = 0.125 * 7.6

A = 0.0967 M

B. ln ( 0.250 / 0.0651) = 0.125 * time

= 10.763 hr

C. T(1/2) = 0.693 / rate constant

= 0.693 / 0.125

= 5.544 hr